OpenStudy (anonymous):

I need help please??? Solve for x 2/3x-2/8=1/4

thomaster (thomaster):

\[\frac{ 2 }{ 3x }-\frac{ 2 }{ 8 }=\frac{ 1 }{ 4 }\]Is that the equation?

OpenStudy (anonymous):

yes!

thomaster (thomaster):

first thing you do is bring -(2/8) to the right. \[\frac{ 1 }{ 4 }+\frac{ 2 }{ 8 } = \frac{ 1 }{ 2 }\]Now the equation will be like\[\frac{ 2 }{ 3x }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

is that all?

thomaster (thomaster):

yep, the answer would be x=1.3333333333333333333...